899:

y=x^3×3√x Please find its derivative

899:

darkknight:

with respect to what variable

899:

Y

darkknight:

dy/dx?

899:

Yes

899:

dy/dx

darkknight:

first you need to use power rule \[x^n = n(x ^{n-1})\]

darkknight:

also you need to use product rule, because you are multiplying 2 functions with a variable (x) term

899:

You can see the question here

darkknight:

yes I am helping you...

899:

I appreciate that kindly do solve it for me please thank you

darkknight:

I am helping you solve it, not solving it for you.

darkknight:

do you know power rule?

899:

Yes

darkknight:

\[\sqrt[3]{x}=x ^{1/3}\] You know power rule, I just told you that. Do you know product rule? (last rule you need to know for this problem)

darkknight:

\[f(x)\times g(x)= f(x) \times g'(x)+g(x) \times f'(x)\] just in case you don't know this is product rule

darkknight:

lets treat x^3 as f(x) and x^(1/3) as another function now we need to apply power rule to both to find the derivatives (of each individually), then apply product rule to find the derivative

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) darkknight first you need to use power rule \[x^n = n(x ^{n-1})\] \(\color{#0cbb34}{\text{End of Quote}}\) following this, x^3 = 3x^(3-1) = 3x^2 correct?

darkknight:

sorry x^3 doesn't equal 3x^2, 3x^2 is the derivative of x^3 (fyi)

darkknight:

applying the same logic, x^(1/3) the derivative is (1/3)(x)^(1/3-1) = \[1/3x ^{-2/3}\]

darkknight:

now you know \[f(x), f'(x), g(x), g'(x)\] apply product rule

899:

Thank You

darkknight:

np : ) have a good day~

899:

899:

Found the answer

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) darkknight \[f(x)\times g(x)= f(x) \times g'(x)+g(x) \times f'(x)\] just in case you don't know this is product rule \(\color{#0cbb34}{\text{End of Quote}}\) f(x) = x^3, f'(x) = 3x^2, g(x) = x^(1/3), g'(x) = 1/3)(x)^(-2/3) If you plug all of these into product rule formula you should get the correct answer

jessiegirl:

1 attachment